I HAVE A QUESTION ON WRITING EQUATIONS OF A LINE THAT PASSES THROUGH THE GIVEN POINT AND IS PERPENDICULAR TO THE GIVEN LINE.
HERE IS THE PROBLEM:
WRITE AN EQUATION OF A LINE THAT PASSES THROUGH THE GIVEN POINT AND IS PERPENDICULAR TO THE GIVEN LINE
(1,-1) Y= 3X+2
HOW DO I DO IT AND PLEASE EXPLAIN FOR ME TO UNDERSTAND THE STEPS

Answer by stanbon(75887)

Write an equation of the line that passes through the given point and is perpendicular to the given line.
(1,-1), y=3x+2
------------
Since the given slope is 3, the slope you want is -1/3.
Equation form is y = mx +b
Substitute to solve for "b":
-1 = (-1/3)*1 + b
b = -1 + (1/3) = -2/3
------------------------
Equation:
y = (-1/3)x + (-2/3)
======================
Cheers,
Stan H.